A) Nail polish remover containing acetone was spilled in a room 8.2 m × 5.0 m × 2.7 m. Measurements indicated that 5,500 mg of acetone evaporated. Calculate the acetone concentration in micrograms per cubic meter.

B) For the 1−hour average value, a city reports 0.14 ppm ozone.

The permissible  limit is 0.12 ppm  ozone. By what  amount, in ppm, does the reported value exceed the permissible limit? This exceeded amount is what percentage of the permissible level? answer in PPM and %
Answer #1

(A) we need to convert 5500 mg to micrograms. In order to that, we need to multiply 5500 by 1000 thus,

5500 mg acetone = 5500 X 1000 = 5500000 micrograms

The volume of the room = length X breadth X height = 8.2 X 5.0 X 2.7 = 110.7 cubic metres

Concentration in micrograms per cubic meter = 5500000/110.7 = 49683.83 micrograms/cubic meter

(B) Permissible limit = 0.12 ppm, Reported average value = 0.14 ppm

The reported limit exceeds the permissible limit.

Difference = 0.14 - 0.12 = 0.02

Percentage = (0.02/0.12)X100 = 16.67 %